OpenStudy (anonymous):

A point P is moving along the curve whose equation is y=((x^3)+17)^(1/2) When P is at (2,5), y is increasing at the rate of 2units/s. How fast is x changing?

OpenStudy (amistre64):

i would simply take the derivative in an implicit manner \[y=(x^3+17)^{1/2}\] \[y'=\frac{2x^2x'}{2}(x^3+17)^{-1/2}\] we know y' so lets solve for x' \[y'=x^2x'(x^3+17)^{-1/2}\] \[y'x^{-2}(x^3+17)^{1/2}=x'\] \[2(2)^{-2}(2^3+17)^{1/2}=x'\] \[2^{-1}(2^3+17)^{1/2}=x'\]